OpenStudy (anonymous):

simplify and write in a+bi form (5-4i)^2 is 25-16i correct?

hartnn (hartnn):

not actually did you use this ? \((x-y)^2 = x^2-2xy+y^2\) ?

OpenStudy (anonymous):

nope. I'll try it (:

OpenStudy (anonymous):

is 5 x and 4y?

hartnn (hartnn):

you mean x=5 and y= 4i then yes

hartnn (hartnn):

\(\Large 5^2 - 2\times 5\times 4i +(4i)^2 =...\) try to simplify :)

OpenStudy (anonymous):

476i?

hartnn (hartnn):

lol how? :P ok, 5^2 = 25 \(\large (4i)^2 = 4^2i^2 = 16i^2=...?\) you know the value of \(\large i^2 \) ??

OpenStudy (anonymous):

115(16i^2)?

hartnn (hartnn):

you know what 'i' is ?

OpenStudy (anonymous):

imaginary

hartnn (hartnn):

did you know \(\huge i=\sqrt{-1}\) ?

hartnn (hartnn):

now you know, so can you tell me what will be \(\large i^2 = (\sqrt{-1})^2=... ?\)

OpenStudy (anonymous):

oh (: 1

hartnn (hartnn):

actually \(\large i^2 = (\sqrt{-1})^2=-1\) :)

OpenStudy (anonymous):

5^2-2x5x4i+16i^2= 15x4+-16= 44?

hartnn (hartnn):

\(\Large 5^2 - 2\times 5\times 4i +(4i)^2 =25 -40i +16i^2\) but \(i^2 = -1\)

hartnn (hartnn):

\(\Large 5^2 - 2\times 5\times 4i +(4i)^2 =25 -40i +16i^2 \\ \large = 25 -40i+16(-1)\) did u get this ? try to simplify further :)

OpenStudy (anonymous):

9-40i?

hartnn (hartnn):

now that, my dear, is absolutely correct! :)

OpenStudy (anonymous):

Thank you beautiful soul :D

hartnn (hartnn):

lol, welcome beautiful.....everything :P ^_^